# GoalLadder: Incremental Goal Discovery with Vision-Language Models

### v1

@@ Abstract
Natural language can offer a concise and human-interpretable means of specifying reinforcement learning (RL) tasks. The ability to extract rewards from a language instruction can enable the development of robotic systems that can learn from human guidance; however, it remains a challenging problem, especially in visual environments. Existing approaches that employ large, pretrained language models either rely on non-visual environment representations, require prohibitively large amounts of feedback, or generate noisy, ill-shaped reward functions. In this paper, we propose a novel method, GoalLadder, that leverages vision-language models (VLMs) to train RL agents from a single language instruction in visual environments. GoalLadder works by incrementally discovering states that bring the agent closer to completing a task specified in natural language. To do so, it queries a VLM to identify states that represent an improvement in agent’s task progress and to rank them using pairwise comparisons. Unlike prior work, GoalLadder does not trust VLM’s feedback completely; instead, it uses it to rank potential goal states using an ELO-based rating system, thus reducing the detrimental effects of noisy VLM feedback. Over the course of training, the agent is tasked with minimising the distance to the top-ranked goal in a learned embedding space, which is trained on unlabelled visual data. This key feature allows us to bypass the need for abundant and accurate feedback typically required to train a well-shaped reward function. We demonstrate that GoalLadder outperforms existing related methods on classic control and robotic manipulation environments with the average final success rate of $\sim$95compared to only $\sim$45of the best competitor.
@@ 1 Introduction
Indeed, LLMs can perform well on open-ended question answering, code generation, and be vast reservoirs of common-sense knowledge [1, 2, 3, 4]. Moreover, multimodal models have even broader applicability to tasks involving images or other modalities [5, 6, 7, 8, 9, 10]. These properties make pretrained language models potentially suitable for providing learning signals to reinforcement learning algorithms, such as by generating preference-based feedback [11] or outputting a reward function directly [12].
Prior work attempts to use LLMs to generate reward functions given access to environment code or interpretable state representations [13, 14, 12], limiting their applicability. Another line of work uses vision-language models (VLMs) to define or learn a reward function given a language instruction and visual observations. In particular, two broad strategies have emerged: (i) embedding-based [2, 15, 16, 17], which aligns visual observations with a language instruction in the embedding space of a VLM; and (ii) preference-based [11], which prompts a VLM to rank segments of an agent’s behaviour by how well they match a textual specification, and trains a reward function from the collected preference data.
Both approaches present certain challenges. Embedding-based methods employing CLIP [5] tend to yield noisy reward functions due to the mismatch between the CLIP training data and the observations from the tested environments [16]. Similarly, these methods require that every observation is embedded to produce rewards, making them expensive and inefficient in using large pretrained models. Preference-based methods like RL-VLM-F [11] produce reward functions with less noise and better correlation with task progress but are nevertheless substantially noisy due to mistakes made by a VLM when comparing trajectories. Erroneous feedback plagues the preference label dataset (at an unknown frequency) making it challenging to robustly train a well-shaped reward function without overfitting to mislabelled samples. Furthermore, although preference-based approaches are more sample-efficient than embedding-based ones, they nevertheless require many VLM queries to train a reward function that generalises well.
@@
To this end, we propose GoalLadder, an algorithm designed to address both of these challenges. Our method leverages a VLM to incrementally discover environment states that bring the agent closer to completing a task. To avoid the problems of prior methods, GoalLadder performs repeated comparisons of a small subset of visual observations and maintains a persistent, ELO-based utility measure (rating) over states. This process allows GoalLadder to progressively refine its estimates of state utilities without being severely affected by noisy VLM feedback. Furthermore, compared to previous methods, GoalLadder requires substantially fewer VLM queries to learn desired behaviours, since rewards are defined as distances between visual observations in a learned embedding space that is trained on unlabelled data, allowing for reward generalisation to unseen states without explicit feedback. Using two classic control and five robotic manipulation tasks, we demonstrate that our method significantly outperforms prior work that utilises vision-language models for RL, with the average final success rate of $\sim$95compared to only $\sim$45of the best competitor. GoalLadder exhibits impressive performance even against the oracle agent that has access to ground-truth reward – nearly matching it across all tested tasks and convincingly surpassing it on one.
@@ 2 Related Work
Recent advances in large language models [18, 19, 20] demonstrate impressive abilities for reasoning and common sense in text and other modalities such as vision [5, 6, 7, 8, 9, 10]. Vision-language models are trained on a joint visual and textual representation space, allowing the capabilities of LLMs to be applied within the visual domain. Nevertheless, existing VLMs suffer from poor spatial understanding [21], making it difficult to reliably apply them to spatial reasoning tasks. Our approach takes this into account by design, using a robust pairwise evaluation strategy to reduce the effects of erroneous feedback on spatial tasks.
@@
The use of language models in reinforcement learning has been predominantly about making LLMs write code, particularly to design a reward function [13, 14, 12]. Although these approaches may improve as language models get better at code comprehension, it is unclear how they would perform in arbitrarily complex physical simulations or transfer to real-world environments. LLMs have also been used to provide a reward signal in text-based environments [22, 23].
@@
Several works explore the integration of VLMs in RL. One avenue is using the embeddings of the VLMs directly in the definition of the reward. For example, [16] use the CLIP model [5] to embed the task description specified in text, as well as an environment’s visual observations, and define the reward as the cosine similarity between the image and text embedding. However, the resulting reward functions tend to be noisy, which is often attributed to the mismatch between the training data of the VLMs and the environments they are applied to [16]. To this end, Baumli et al. [17] fine-tune a CLIP model with contrastive learning to output the success of an agent in reaching a text-based goal. Other papers use vision-language models to align visual observations with language descriptions of a task [2, 24, 15, 25], to perform incremental trajectory improvement from human feedback [26] or to elicit better task-relevant representations [27].
VLMs can also be used for direct feedback to learn a reward function from VLM goal-conditioned preferences [11]. In particular, Wang et al. [11] use VLMs to provide preference labels over states of visual observations conditioned on a task description and use these labels to learn a reward function with Reinforcement Learning from Human Preferences (RLHF) [28]. Although this work improves upon previous methods, noise in the learned reward function remains evident. Furthermore, for the experiments with robotic manipulation, the authors remove the robot from the image to improve the VLM’s ability to identify goal states in object-oriented tasks. This trick, however, also reduces the amount of information a VLM can use to identify incrementally better states. In contrast, GoalLadder solves the tasks without environment modifications and utilises an ELO-based rating system to reduce the effects of noisy VLM feedback.
@@ 3.1 Preliminaries
We formulate the problem as a Markov Decision Process (MDP), defined as a tuple =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$. Here, $\mathcal{S}$ represents the state space of the environment, $\mathcal{A}$ is the agent’s action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ defines the transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. At each timestep $t$, an agent takes an action $a_{t}$, transitions to state $s_{t+1}$, and receives reward $r_{t}$. In our setup, we assume that the agent similarly receives an image observation $o_{t}$ that represents the state of the environment. The goal of the agent is to maximise the sum of discounted rewards in the environment, defined as $G=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$.
@@
We use soft-actor critic (SAC) [29] as the RL backbone of our agent. However, in GoalLadder, the reward function changes periodically as new targets are being set over the course of training. The vanilla implementation of SAC can struggle with non-stationary reward given that it reuses past data from the replay buffer, which may not be up to date with respect to the latest reward function. As such, similar to [30], we periodically relabel all of the agent’s trajectories with the latest reward function.
@@ 3.2.1 Discovery of candidate goals
where $y=-1$ indicates that there is no decision or the images are equivalent, $y=0$ means $g^{*}$ is deemed a better goal than $o_{j}$, and $y=1$ means $o_{j}$ is deemed a better goal than $g^{*}$.
If $y=1$, the newly sampled state $o_{j}$ is inserted into $\mathcal{B}_{g}$ as a fresh candidate goal with a standard initial rating. Conversely, if $y=0$ or $y=-1$, $o_{j}$ is disregarded in order to keep the goal buffer focused on higher-quality candidates.
@@ 3.2.2 Rating of candidate goals
Ideally, our goal ratings would have two main properties: (1) be robust to noisy VLM outputs and (2) allow for quick updates to a candidate goal’s rating when new and better goals are discovered. In GoalLadder, we draw inspiration from the ELO chess rating system [31] to maintain a robust, adaptive measure of each candidate goal’s utility. In chess, players’ ratings are updated after every match based on two factors: the observed outcome (win, lose, or draw) and the expected outcome given the players’ current ratings.
@@ 3.2.3 Defining the reward function
To address this, GoalLadder learns a visual feature extractor to produce a compact latent representation of environment observations. We achieve this using a variational autoencoder training objective [32],
@@
By training on a diverse set of observations the agent encounters, we obtain a latent space that captures salient visual (and potentially semantic) features of the environment. Finally, we define the agent’s reward at time step $t-1$ to be:
@@
As the candidate goals are being discovered and rated during training, we periodically update our reward function and relabel all stored transitions [30] with respect to the top-rated candidate goal. Periodicity ensures that sudden or erroneous changes in the top-rated goal do not significantly destabilise the training process, while also keeping the best goal estimate up-to-date.
@@ 4 Experiments
We run a variety of continuous-control experiments to investigate GoalLadder’s performance. In particular, we use two classic control environments (CartPole, MountainCar) from OpenAI Gym [33] and five robotic manipulation environments (Drawer Close, Drawer Open, Sweep Into, Window Open, Button Press) from the Metaworld suite [34]. Importantly, unlike prior works [16, 11], we do not perform any special environment modifications to help VLMs discern task progress. We use feedback rates of $K=2000$ with $M=5$ for OpenAI Gym environments and $K=500$ with $M=5$ for the Metaworld environments.
@@
  • VLM-RM [16]: Uses CLIP [5] to embed the task description and an image observation and defines the reward as the cosine-similarity between the two embeddings.
  • RoboCLIP [15]: Employs a pre-trained video-language model, S3D [35], to compute rewards based on the similarity between an agent’s video trajectory and either a reference video or a text description. To ensure a fair comparison, we use the text version of RoboCLIP.
  • RL-VLM-F [11]. Uses a VLM to compare pairs of images with respect to a language instruction. Unlike GoalLadder, RL-VLM-F utilises VLM feedback to train a reward function from the collected labels. For this baseline, we apply the same feedback rate used in GoalLadder and use the same vision-language model as backbone – Gemini 2.0 Flash.
@@ 4.2 Can GoalLadder solve the tasks?
Figure 4 shows that GoalLadder can effectively solve continuous control tasks given only a language instruction, reaching a mean success rate of $\sim$95averaged across all tasks. Curiously, our method nearly matches the performance of the ground-truth reward baseline and even convincingly surpasses it for the Drawer Open task. Empirically, we find that sometimes the oracle agent learns to reach the drawer, but does not learn to pull the handle to open it. We believe this once again highlights the difficulty of manually designing effective reward functions – an issue that was also previously discussed in Wang et al. [11].
The preference-based baseline, RL-VLM-F, manages to solve the relatively easy tasks – Cartpole and MountainCar from the OpenAI Gym suite, as well as Drawer Close from Metaworld. However, it begins to struggle with the more complicated Metaworld tasks. We attribute this poor performance of RL-VLM-F to the fact that it requires more feedback to learn an effective reward function, as well as ways to mitigate the effects of noisy labels in the preference data. Similarly, VLM-RM and RoboCLIP struggle to learn in most of the tasks, which is in line with previously observed limitations of these methods [11].
Figure 4: GoalLadder performance against baselines on OpenAI Gym and Metaworld environments. Shaded regions represent the standard deviation over 3 seeds. Averaged across all tasks, GoalLadder achieves a mean success rate of $\sim$95%, compared to the next best competitor performance of $\sim$45achieved by RL-VLM-F.

### v2

@@ Abstract
Natural language can offer a concise and human-interpretable means of specifying reinforcement learning (RL) tasks. The ability to extract rewards from a language instruction can enable the development of robotic systems that can learn from human guidance; however, it remains a challenging problem, especially in visual environments. Existing approaches that employ large, pretrained language models either rely on non‑visual environment representations, require prohibitively large amounts of feedback, or generate noisy, ill‑shaped reward functions. In this paper, we propose a novel method, GoalLadder, that leverages vision-language models (VLMs) to train RL agents from a single language instruction in visual environments. GoalLadder works by incrementally discovering states that bring the agent closer to completing a task specified in natural language. To do so, it queries a VLM to identify states that represent an improvement in agent’s task progress and to rank them using pairwise comparisons. Unlike prior work, GoalLadder does not trust VLM’s feedback completely; instead, it uses it to rank potential goal states using an ELO-based rating system, thus reducing the detrimental effects of noisy VLM feedback. Over the course of training, the agent is tasked with minimising the distance to the top-ranked goal in a learned embedding space, which is trained on unlabelled visual data. This key feature allows us to bypass the need for abundant and accurate feedback typically required to train a well-shaped reward function. We demonstrate that GoalLadder outperforms existing related methods on classic control and robotic manipulation environments with the average final success rate of $\sim$95% compared to only $\sim$45% of the best competitor.
@@ 1 Introduction
Indeed, LLMs can perform well on open-ended question answering, code generation, and be vast reservoirs of common-sense knowledge [he_can_2024, cui_can_2022, petroni_language_2019, kritharoula_language_2023]. Moreover, multimodal models have even broader applicability to tasks involving images or other modalities [radford_learning_2021, jia_scaling_2021, yao_filip_2021, zhou_learning_2022, gao_clip-adapter_2024, bordes_introduction_2024]. These properties make pretrained language models potentially suitable for providing learning signals to reinforcement learning algorithms, such as by generating preference-based feedback [wang_rl-vlm-f_2024] or outputting a reward function directly [ma_eureka_2024].
Prior work attempts to use LLMs to generate reward functions given access to environment code or interpretable state representations [yu_language_2023, xie_text2reward_2024, ma_eureka_2024], limiting their applicability. Another line of work uses vision-language models (VLMs) to define or learn a reward function given a language instruction and visual observations. In particular, two broad strategies have emerged: (i) embedding-based [cui_can_2022, sontakke_roboclip_2023, rocamonde_vision-language_2023, baumli_vision-language_2024], which aligns visual observations with a language instruction in the embedding space of a VLM; and (ii) preference-based [wang_rl-vlm-f_2024], which prompts a VLM to rank segments of an agent’s behaviour by how well they match a textual specification, and trains a reward function from the collected preference data.
Both approaches present certain challenges. Embedding-based methods employing CLIP [radford_learning_2021] tend to yield noisy reward functions due to the mismatch between the CLIP training data and the observations from the tested environments [rocamonde_vision-language_2023]. Similarly, these methods require that every observation is embedded to produce rewards, making them expensive and inefficient in using large pretrained models. Preference-based methods like RL-VLM-F [wang_rl-vlm-f_2024] produce reward functions with less noise and better correlation with task progress but are nevertheless substantially noisy due to mistakes made by a VLM when comparing trajectories. Erroneous feedback plagues the preference label dataset (at an unknown frequency) making it challenging to robustly train a well-shaped reward function without overfitting to mislabelled samples. Furthermore, although preference-based approaches are more sample-efficient than embedding-based ones, they nevertheless require many VLM queries to train a reward function that generalises well.
@@
To this end, we propose GoalLadder, an algorithm designed to address both of these challenges. Our method leverages a VLM to incrementally discover environment states that bring the agent closer to completing a task. To avoid the problems of prior methods, GoalLadder performs repeated comparisons of a small subset of visual observations and maintains a persistent, ELO-based utility measure (rating) over states. This process allows GoalLadder to progressively refine its estimates of state utilities without being severely affected by noisy VLM feedback. Furthermore, compared to previous methods, GoalLadder requires substantially fewer VLM queries to learn desired behaviours, since rewards are defined as distances between visual observations in a learned embedding space that is trained on unlabelled data, allowing for reward generalisation to unseen states without explicit feedback. Using two classic control and five robotic manipulation tasks, we demonstrate that our method significantly outperforms prior work that utilises vision-language models for RL, with the average final success rate of $\sim$95% compared to only $\sim$45% of the best competitor. GoalLadder exhibits impressive performance even against the oracle agent that has access to ground-truth reward – nearly matching it across all tested tasks and convincingly surpassing it on one.
@@ 2 Related Work
Recent advances in large language models [chowdhery_palm_2023, touvron_llama_2023, openai_gpt-4_2024] demonstrate impressive abilities for reasoning and common sense in text and other modalities such as vision [radford_learning_2021, jia_scaling_2021, yao_filip_2021, zhou_learning_2022, gao_clip-adapter_2024, bordes_introduction_2024]. Vision-language models are trained on a joint visual and textual representation space, allowing the capabilities of LLMs to be applied within the visual domain. Nevertheless, existing VLMs suffer from poor spatial understanding [kamath_whats_2023], making it difficult to reliably apply them to spatial reasoning tasks. Our approach takes this into account by design, using a robust pairwise evaluation strategy to reduce the effects of erroneous feedback on spatial tasks.
@@
The use of language models in reinforcement learning has been predominantly about making LLMs write code, particularly to design a reward function [yu_language_2023, xie_text2reward_2024, ma_eureka_2024]. Although these approaches may improve as language models get better at code comprehension, it is unclear how they would perform in arbitrarily complex physical simulations or transfer to real-world environments. LLMs have also been used to provide a reward signal in text-based environments [klissarov_motif_2023, kwon_reward_2023].
@@
Several works explore the integration of VLMs in RL. One avenue is using the embeddings of the VLMs directly in the definition of the reward. For example, [rocamonde_vision-language_2023] use the CLIP model [radford_learning_2021] to embed the task description specified in text, as well as an environment’s visual observations, and define the reward as the cosine similarity between the image and text embedding. However, the resulting reward functions tend to be noisy, which is often attributed to the mismatch between the training data of the VLMs and the environments they are applied to [rocamonde_vision-language_2023]. To this end, baumli_vision-language_2024 fine-tune a CLIP model with contrastive learning to output the success of an agent in reaching a text-based goal. Other papers use vision-language models to align visual observations with language descriptions of a task [cui_can_2022, fan_minedojo_2022, sontakke_roboclip_2023, nam_lift_2023], to perform incremental trajectory improvement from human feedback [yang_trajectory_2024] or to elicit better task-relevant representations [chen_vision-language_2024].
VLMs can also be used for direct feedback to learn a reward function from VLM goal-conditioned preferences [wang_rl-vlm-f_2024]. In particular, wang_rl-vlm-f_2024 use VLMs to provide preference labels over states of visual observations conditioned on a task description and use these labels to learn a reward function with Reinforcement Learning from Human Preferences (RLHF) [christiano_deep_2017]. Although this work improves upon previous methods, noise in the learned reward function remains evident. Furthermore, for the experiments with robotic manipulation, the authors remove the robot from the image to improve the VLM’s ability to identify goal states in object-oriented tasks. This trick, however, also reduces the amount of information a VLM can use to identify incrementally better states. In contrast, GoalLadder solves the tasks without environment modifications and utilises an ELO-based rating system to reduce the effects of noisy VLM feedback.
@@ 3.1 Preliminaries
We formulate the problem as a Markov Decision Process (MDP), defined as a tuple $\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$. Here, $\mathcal{S}$ represents the state space of the environment, $\mathcal{A}$ is the agent’s action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ defines the transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. At each timestep $t$, an agent takes an action $a_{t}$, transitions to state $s_{t+1}$, and receives reward $r_{t}$. In our setup, we assume that the agent similarly receives an image observation $o_{t}$ that represents the state of the environment. The goal of the agent is to maximise the sum of discounted rewards in the environment, defined as $G=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$.
@@
We use soft-actor critic (SAC) [haarnoja_soft_2018] as the RL backbone of our agent. However, in GoalLadder, the reward function changes periodically as new targets are being set over the course of training. The vanilla implementation of SAC can struggle with non-stationary reward given that it reuses past data from the replay buffer, which may not be up to date with respect to the latest reward function. As such, similar to [lee_pebble_2021], we periodically relabel all of the agent’s trajectories with the latest reward function.
@@ 3.2.1 Discovery of candidate goals
where y=1y=-1 indicates that there is no decision or the images are equivalent, $y=0$ means $g^{*}$ is deemed a better goal than $o_{j}$, and $y=1$ means $o_{j}$ is deemed a better goal than $g^{*}$.
If $y=1$, the newly sampled state $o_{j}$ is inserted into $\mathcal{B}_{g}$ as a fresh candidate goal with a standard initial rating. Conversely, if $y=0$ or y=1y=-1, $o_{j}$ is disregarded in order to keep the goal buffer focused on higher-quality candidates.
@@ 3.2.2 Rating of candidate goals
Ideally, our goal ratings would have two main properties: (1) be robust to noisy VLM outputs and (2) allow for quick updates to a candidate goal’s rating when new and better goals are discovered. In GoalLadder, we draw inspiration from the ELO chess rating system [elo_rating_2008] to maintain a robust, adaptive measure of each candidate goal’s utility. In chess, players’ ratings are updated after every match based on two factors: the observed outcome (win, lose, or draw) and the expected outcome given the players’ current ratings.
@@ 3.2.3 Defining the reward function
To address this, GoalLadder learns a visual feature extractor to produce a compact latent representation of environment observations. We achieve this using a variational autoencoder training objective [kingma_auto-encoding_2014],
@@
By training on a diverse set of observations the agent encounters, we obtain a latent space that captures salient visual (and potentially semantic) features of the environment. Finally, we define the agent’s reward at time step t1t-1 to be:
@@
As the candidate goals are being discovered and rated during training, we periodically update our reward function and relabel all stored transitions [lee_pebble_2021] with respect to the top-rated candidate goal. Periodicity ensures that sudden or erroneous changes in the top-rated goal do not significantly destabilise the training process, while also keeping the best goal estimate up-to-date.
@@ 4 Experiments
We run a variety of continuous-control experiments to investigate GoalLadder’s performance. In particular, we use two classic control environments (CartPole, MountainCar) from OpenAI Gym [brockman_openai_2016] and five robotic manipulation environments (Drawer Close, Drawer Open, Sweep Into, Window Open, Button Press) from the Metaworld suite [yu_meta-world_2021]. Importantly, unlike prior works [rocamonde_vision-language_2023, wang_rl-vlm-f_2024], we do not perform any special environment modifications to help VLMs discern task progress. We use feedback rates of $K=2000$ with $M=5$ for OpenAI Gym environments and $K=500$ with $M=5$ for the Metaworld environments.
@@
  • VLM-RM [rocamonde_vision-language_2023]: Uses CLIP [radford_learning_2021] to embed the task description and an image observation and defines the reward as the cosine-similarity between the two embeddings.
  • RoboCLIP [sontakke_roboclip_2023]: Employs a pre-trained video-language model, S3D [zhong_s3d_2024], to compute rewards based on the similarity between an agent’s video trajectory and either a reference video or a text description. To ensure a fair comparison, we use the text version of RoboCLIP.
  • RL-VLM-F [wang_rl-vlm-f_2024]. Uses a VLM to compare pairs of images with respect to a language instruction. Unlike GoalLadder, RL-VLM-F utilises VLM feedback to train a reward function from the collected labels. For this baseline, we apply the same feedback rate used in GoalLadder and use the same vision-language model as backbone – Gemini 2.0 Flash.
@@ 4.2 Can GoalLadder solve the tasks?
Figure 4 shows that GoalLadder can effectively solve continuous control tasks given only a language instruction, reaching a mean success rate of $\sim$95% averaged across all tasks. Curiously, our method nearly matches the performance of the ground-truth reward baseline and even convincingly surpasses it for the Drawer Open task. Empirically, we find that sometimes the oracle agent learns to reach the drawer, but does not learn to pull the handle to open it. We believe this once again highlights the difficulty of manually designing effective reward functions – an issue that was also previously discussed in wang_rl-vlm-f_2024.
The preference-based baseline, RL-VLM-F, manages to solve the relatively easy tasks – Cartpole and MountainCar from the OpenAI Gym suite, as well as Drawer Close from Metaworld. However, it begins to struggle with the more complicated Metaworld tasks. We attribute this poor performance of RL-VLM-F to the fact that it requires more feedback to learn an effective reward function, as well as ways to mitigate the effects of noisy labels in the preference data. Similarly, VLM-RM and RoboCLIP struggle to learn in most of the tasks, which is in line with previously observed limitations of these methods [wang_rl-vlm-f_2024].
Figure 4: GoalLadder performance against baselines on OpenAI Gym and Metaworld environments. Shaded regions represent the standard deviation over 3 seeds. Averaged across all tasks, GoalLadder achieves a mean success rate of $\sim$95%, compared to the next best competitor performance of $\sim$45% achieved by RL-VLM-F.